OpenStudy (anonymous):

WILL FAN + MEDAL! Solve the system by linear combination. -x+4y= -1 2x-y= 5.5

OpenStudy (anonymous):

-x+4y= -1 ...........(1 2x-y= 5.5 ..........(2 Multiplying eq (1 by 2 and then adding it with eq (2 we find -2x+8y= -2 ...........(1 2x-y= 5.5 ..........(2 --------------------- 7y = 3.5 y= 3.5/7 y= 0.5 Substituting y= 0.5 in eq (1 we find -x+4 *0.5= -1 -x+2= -1 1+2= x x=3 hence (x, y) = (3, 0.5)

OpenStudy (anonymous):

@daisysosa512